OpenStudy (unklerhaukus):

what volume in \([\text{mm}^3]\) is one grain of \(\text{Pt}\) \[1[\text{gr}]\text{Pt}=\]

OpenStudy (unklerhaukus):

\[7000[\text{gr}]=1[\text {lb}]=453.59[\text g]\] \[1[\text{gr}]=\frac{453.59}{7000}[\text g]\approx 0.064798[\text g]\]

OpenStudy (unklerhaukus):

\[M_\text{Pt}=195.078[\text {g/mol}]\]

OpenStudy (unklerhaukus):

\[1[\text{gr}]\text{Pt}=3.3217\times10^{-4}[\text{mole}]\text{Pt}\]

OpenStudy (unklerhaukus):

How do i get to volume ?

OpenStudy (unklerhaukus):

i know how to convert volumes , but i dont know the volume of a mole of platinum

OpenStudy (unklerhaukus):

i guess i just need the density

OpenStudy (unklerhaukus):

\[\rho_{\text{Pt}}=\frac{m_{\text{Pt}}}{V_{\text{Pt}}}\] \[{V_{\text{Pt}}}=\frac{m_{\text{Pt}}}\rho_{\text{Pt}}\]

OpenStudy (unklerhaukus):

\[m_\text{Pt}=1[\text{gr}]=0.064798[\text g]\] \[\rho_{\text{Pt}}=21.45 [\text{g/cm}]^3\] \[{V_{\text{Pt}}}=\frac{m_{\text{Pt}}}\rho_{\text{Pt}}=\frac{0.064798[\text g]}{21.45 [\text{g/cm}]^3}=0.0030209[\text{cm}]^3\]

OpenStudy (unklerhaukus):

\[{V_{\text{Pt}}}=0.0030209[\text{cm}]^3\times\left(\frac{10[\text {mm}]}{[\text{cm}]}\right)^3\] \[ V_{\text{Pt}}=3[\text{mm}]^3\]

OpenStudy (unklerhaukus):

is this right?

OpenStudy (unklerhaukus):

\[7000[\text{gr}]=16[\text{oz}]=1[\text {lb}]\]

OpenStudy (unklerhaukus):

\[\frac 1{437.5} [\text{oz}] \text{Pt}\]

OpenStudy (unklerhaukus):

\[1[\text{gr}]\text{Pt}\approx3[\text{mm}]\color\red \checkmark\]

OpenStudy (unklerhaukus):

1/500 =0.002000000

OpenStudy (unklerhaukus):